C&A produces three kinds of bread: wheat, white, and multigrain. The data on these products are given below:

 

Wheat

White

Multigrain

Demand (lbs per hour)

50

20

10

Changeover time (min)

60

30

45

Production rate (lbs per hour)

200

200

200

How many lbs of white bread should C&A produce with each batch in order to minimize inventory while satisfying demand?

O 37.5

O 75

O 187.5

O 300
Answer #1

Total change over time = 60 + 30 + 45 = 135 minutes = 2.25 hours

Total Demand = 50 + 20 + 10 = 80 lbs / hr

Processing time for Wheat bread = (1 / 200) hr

Processing time for White bread = (1 / 200) hr

Processing time for Multigrain bread = (1 / 200) hr

Let the batch size be Z

Total Processing time = (1/200) * Z

Total Batch time = Total processing time + Total Set-up time = (1/200)*Z + 2.25

Process capacity = Total output / Total time = Z / ((1/200)*Z + 2.25)

Capacity required = Total Demand = 80 lbs / hr

Hence, Z / ((Z/200) + 2.25) = 80

Z = 80 * ((Z / 200 ) + 2.25)

Z * 200 = 80Z + 80*200*2.25

Z = 300 lbs

Total batch size = 300 lbs

Proportion of White bread = (50/80) * 300 = 75 lbs

Answer: 187.5

-----------------------------------------------------------------------------------------------------------------------

In case of any doubt, please ask through the comment section before Upvote/downvote.

If you liked the answer, please give an upvote. This will be quite encouraging for me, thank-you!!!